OpenStudy (anonymous):

help me simplify this plynomial please: sqrt x^5 y^7/ x^33 y^1/4 asap plz

OpenStudy (anonymous):

is it the square root of the whole thing? \[\sqrt{\frac{x^5y^7}{x^{33}y^{\frac{1}{4}}}}\]

OpenStudy (anonymous):

srry only the 1st is sqrt nd x^3 not 33 my bad

OpenStudy (anonymous):

sqrt x^5y^7

OpenStudy (anonymous):

\[\frac{\sqrt{x^5y^7}}{x^3y^{\frac{1}{4}}}\]

OpenStudy (anonymous):

like that?

OpenStudy (anonymous):

yeaa :)

OpenStudy (anonymous):

ok we go quickly \[\frac{(x^5y^7)^{\frac{1}{2}}}{x^3y^{\frac{1}{4}}}\]

OpenStudy (anonymous):

\[\frac{x^{\frac{5}{2}}y^{\frac{7}{2}}}{x^3y^{\frac{1}{4}}}\]

OpenStudy (anonymous):

\[3-\frac{5}{2}=\frac{1}{2}\] \[\frac{7}{2}-\frac{1}{4}=\frac{13}{4}\] so you get \[\frac{y^{\frac{13}{4}}}{x^{\frac{1}{2}}}\]

OpenStudy (anonymous):

you can write this as \[\frac{y^3\sqrt[4]{y}}{\sqrt{x}}\] if you like

OpenStudy (anonymous):

thaaank k you soo much